OpenStudy (anonymous):

Find the length of the radius of the given circle. (x - 5)2 + (y + 3)2 = 25 5 12.5 25 Please help

OpenStudy (anonymous):

B

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

Let's solve for x. (x−5)(2)+(y+3)(2)=25 Step 1: Add 4 to both sides. 2x+2y−4+4=25+4 2x+2y=29 Step 2: Add -2y to both sides. 2x+2y+−2y=29+−2y 2x=−2y+29 Step 3: Divide both sides by 2. 2x2=−2y+292 x=−y+292 Answer: x=−y+29/2 29/2=14.5 14.5/2= 7.25 7.25/2=12.5

OpenStudy (anonymous):

@karatechopper is that correct??

OpenStudy (anonymous):

want me to explain it better?

OpenStudy (anonymous):

2x/2=−2y+29/2 Let's solve for x. (x−5)(2)+(y+3)(2)=25 Step 1: Add 4 to both sides. 2x+2y−4+4=25+4 2x+2y=29 Step 2: Add -2y to both sides. 2x+2y+−2y=29+−2y 2x=−2y+29 Step 3: Divide both sides by 2. 2x/2=−2y+29/2 x=−y+292 SIMPLIFY: 29/2=14.5 14.5/2= 7.25 7.25/2= Answer: 12.5

OpenStudy (anonymous):

that's the best I can explain.

OpenStudy (anonymous):

It is correct.

OpenStudy (karatechopper):

The equation of a circle is, (x−h)^2+(y−k)^2=r^2 So if we look at the given equation of the problem, (x - 5)2 + (y + 3)2 = 25 Which part of the given equation gives you a hint about the radius in reference to out model equation of the circle?

OpenStudy (karatechopper):

@vbroberson I humbly don't agree with your answer.

OpenStudy (anonymous):

@karatechopper How did I get it wrong then?

OpenStudy (karatechopper):

Vbroberson I will dm you. Jc can you answer the question I had above here?

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

I see where I messed up.

OpenStudy (anonymous):

The radius would be 5